6. A 25.00mL sample of an H2SO4 solution of unknown concentration requires 24.16mL of 0.106M sodium hydroxide for complete neutralization (to reach the equivalence point). What was the concentration of the H2SO4 solution?
Answer #1

Balanced chemical reaction equation is:

H2SO4 (aq) + 2NaOH (aq) ==> 2H2O (l) + Na2SO4 (aq)

1 mol acid requires 2 mol base

moles base = 24.16 mL x 10-3 L x 0.106 M = 0.00256096 moles

moles acid required = 0.00256096 / 2 = 0.00128048 moles

volume of acid = 25 mL = 0.025 L

concentration = moles / liter = 0.00128048 moles / 0.025 L = 0.0512192 M

concentration of the H2SO4 solution = 0.05122 M

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")